OpenStudy (anonymous):

(sin x) - 2 (sin x)^2 = 0. How do I solve the equation in the interval [0, 2pi) ?

OpenStudy (anonymous):

factoring sinx( 1 - 2sinx) = 0 sinx = 0 or 1- 2sinx = 0 sinx = 0 => x = 0, Pi, 2pi 1-2sinx = 0 => sinx = 1/2 => x = pi/6, 5pi/ 6

OpenStudy (anonymous):

1) sinx=0... x=0,pi 2)sinx = 1/2 .. x= pi/6, 5pi/6 it can't be 2pi since the interval does not include 2pi

OpenStudy (anonymous):

um yes, Nitin is right...I missed that part...so the answer included, 0, pi, pi/6 and 5pi/6

OpenStudy (anonymous):

wait would it be 1/2 because you're dividing by the -2? or is santas_little_helper right?

OpenStudy (anonymous):

1 - 2sinx = 0 1 = 2sinx => sinx = 1/2

OpenStudy (anonymous):

okay i see now, thanks so much :)

OpenStudy (anonymous):

:) no problem guys